OpenStudy (anonymous):

Throughout the day, the temperature outside of John's house changed from negative eleven degrees Fahrenheit to fifteen degrees Fahrenheit.How much did the temperature outside increase?

OpenStudy (anonymous):

15 - (-11) = 15 + 11 = 26 degrees <== the increase